OpenStudy (anonymous):

mr a can buy six cds and 3 videos for $163.17 or nine cd's and two vids for $200.53.what is the cost of 1 cd?

OpenStudy (callisto):

let x and y be the cost 1CD and 1 video respectively 6x + 3y = 163.17 => y = 54.39 - 2x -(1) 9x +2y = 200.53 -(2) Put (1) into (2) and solve for x. There you'll get your answer. Can you do it?

OpenStudy (anonymous):

6C+3V=163.17 9C+2V=200.53 C=CD V=VIDEO 6C=163.17-3V C=(163.17)/(6)-(3V)/(6) substitute to the second equation.... and you'll have.. 9C+2V=200.53 \[\Large 9\left(\frac{163.17}{6}-\frac{3V}{6}\right)+2V=200.53\] Go on...